OpenStudy (anonymous):

{Please help! I don't understand it. I'll give you Medal+Fan } 1. Is -7+9= -9+7, true, false, or open? a.true b. false c. open

OpenStudy (anonymous):

Is -7+9= -9+7, true, false, or open? 2 = -2 False

OpenStudy (anonymous):

okay, thank you so much ^_^

OpenStudy (anonymous):

I have a few more questions :) please someone

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

2. Is 4x-3= 19 true, false, or open? a. true. b. false c. open help me solve it pretty please :3

OpenStudy (anonymous):

2. Is 4x-3= 19 true, false, or open? 4x-3 = 19 4x-3 + 3 = 19 +3 4x = 22 4x/4 = 22/4 x = 22/4 OPEN